Zeronknight (zeronknight):

Jan, Marsha and Cindy each compared their chore charts. Jan has completed 8 chores and has 3 left to complete. Marsha has completed 9 and has 4 left to complete. Cindy has completed 5 and has 3 more to complete. Who has the highest ratio of completed to not completed chores? A. Marsha B. Jan C. Cindy D. They are all the same

OpenStudy (superteenblue):

sup

OpenStudy (superteenblue):

this is so easy what grade you in

OpenStudy (superteenblue):

marsha

OpenStudy (superteenblue):

i think i willl un fan you zeronight

Zeronknight (zeronknight):

Um.......okay then.

OpenStudy (superteenblue):

i said maybe...... dont you remember i helped you?

OpenStudy (superteenblue):

long ago

Zeronknight (zeronknight):

Well can you make this easy for me? :P

OpenStudy (mathstudent55):

Answer is not Marsha.

OpenStudy (superteenblue):

it is so!

OpenStudy (mathstudent55):

Can you show the work you did and why that is the correct answer?

OpenStudy (mathstudent55):

What are the three ratios that you need to compare?

OpenStudy (superteenblue):

dude....xd...... ok xd......sorry i cant stop laughing so they say witch person has more chores to do and it is marsha because she has 4 and the rest have 3

OpenStudy (mathstudent55):

The question is who has the highest ratio of completed to not completed chores. First, let's look at the ratios. Then we need to compare them. Jan: 8/3 Marsha: 9/4 Cindy: 5/3 Those are the ratios. To compare them quickly, take a calculator and find a decimal for each ratio. Jan: 8/3 = 2.666... Marsha: 9/4 = 2.25 Cindy: 5/3 = 1.666... Now you can compare easily. Which ratio is the greatest number of the three?

Zeronknight (zeronknight):

Oh so when comparing ratios, for example. 5/4 3/5 7/4 I would divide 5 by 4, and so on? And Jan has the greatest number. @mathstudent55

OpenStudy (superteenblue):

oh.....

OpenStudy (superteenblue):

talk to you later!

OpenStudy (mathstudent55):

@Zeronknight Yes, that is correct. Since Jan's ratio is 2.666..., and that is greater than 2.25 and than 1.666..., Jan's ratio is the greatest in this problem.